# Retrieval-Augmented Mixture of LoRA Experts for Uploadable Machine Learning

### v1

@@ 5.0.1. Training And Evaluation Datasets
Table 1. We report the average performance of each task cluster. The full results of each task are shown in Tab.2 & 3. ”IID” signifies that RAMoLE can access any LoRA for every test sample, encompassing the LoRA specific to the sample’s task. ”OOD” indicates that for each test sample, we mask the LoRA associated with its specific task during the retrieval phase. Consequently, no sample can access its ideal LoRA, allowing us to assess the RAMoLE’s cross-task generalization capability. The performance of perfectly selected corresponding LoRA for each sample is colored in gray. We have bolded the best performance of each task and underlined the best performance in the ”OOD” setting.
5.0.3. Mixed Task Evaluation Dataset.
@@
For constructing the mixed-task dataset, we randomly chose 50 samples from the test set for each task used in training 48 LoRAs, subsequently mixing and shuffling these samples to form a unified dataset with 2400 data entries.
@@
Table 2. Mixed Tasks evaluation on both NLU & NLG tasks. “OOD” indicates that during retrieval, we masked the corresponding task’s LoRA for testing generalization when facing unknown tasks.
@@
Table 3. Mixed Tasks evaluation on both NLU & NLG tasks. “OOD” indicates that during retrieval, we masked the corresponding task’s LoRA for testing generalization when facing unknown tasks.
5.0.4. Baseline Methods.
We compared our method with the following baselines: (1) Mixture of Experts  (Zhu et al., 2023; Zadouri et al., 2023; Liu et al., 2023; Wang et al., 2022; Wu et al., 2023a). Many works have considered coordinating different adapters through MoE, and here we explored three distinct variants: one employing a soft mixture of experts and the other utilizing discrete routing (top1 and top3).It is worth noting that in this paper, our primary focus is on post-hoc MoE training. This implies that the LoRA experts have been previously trained, and only the parameters of the router are trainable during the training of the MoE methods. (2) SMEAR  (Muqeeth et al., 2023) introduces the concept of adaptive routing by performing a weighted average of different adapters’ parameters to utilize various experts effectively. For the MoE and SMEAR baselines, challenges arise in scaling due to training confined to a limited set of LoRAs. Consequently, we strategically selected a dedicated LoRA expert for each domain to specialize in router training. (3) AdapterSoup  (Chronopoulou et al., 2023) uniformly selects the corresponding LoRAs for the entire downstream task, which lacks the ability to provide personalized service for diverse requests. (4) LoRAHub  (Huang et al., 2023) enables black-box optimization to learn the weights of various LoRA parameters, thereby facilitating weighted parameter averaging for specific downstream tasks. In our implementation, we conformed to the default setting, which entails randomly selecting 20 LoRAs from the available LoRA pool and performing weighted parameter averaging. For the MoE, SMEAR, and LoRAHub approaches, we selected 20 data samples from the training datasets of all tasks to serve as their training data.
Method Top 1 Top 3 Top 5 Top 8
all-mpnet-base-v2 58.40 78.26 84.77 90.24
all-MiniLM-L6-v2 51.73 73.11 80.54 87.18
msmarco-distilbert-cos-v5 45.84 66.01 75.14 82.67
gtr-t5-xl 53.19 69.72 77.41 83.59
LoraRetriever 0% 60.80 79.29 85.57 91.58
LoraRetriever 40% 63.16 89.09 95.45 98.97
LoraRetriever 100% 74.08 97.37 99.15 99.82
Table 4. Comparison of Sentence Embedding Techniques in LoRA Retrieval: The notation LoraRetrieverksignifies that the model underwent supplementary training on $k$ percent of the tasks. The performance of the selected retriever model in the evaluation phase is highlighted in gray.
5.0.5. Implementation of Baseline Methods.
(16) $y=\sum_{i=i}^{k}R(x)_{i}E_{i}(x).$
We implied two variants of the MoE routing mechanism. (1) Dense Gating. Following  (Zadouri et al., 2023), the router network consists of a dense layer with trainable parameter $W_{g}$, and the gating score could be obtained through a softmax function by:
(17) $s_{i}=R(x)_{i}=softmax(W_{g}^{T}x),$
(2)Sparse Gate. To maintain the sparsity while training, we leverage the Gumbel softmax trick as  (Muqeeth et al., 2023; Nie et al., 2021), where the router can be written as:
(18) $\hat{R}(x)_{i}=\frac{(log(R(x)_{i})+g_{i})/\tau}{\sum_{i=1}^{k}exp((log(R(x)_{i})+g_{i})/\tau)}$
where $g_{i}\sim\text{Gumbel}(0,1)$ and $\tau$ is the temperature.
SMEAR SMEAR  (Muqeeth et al., 2023) does not perform routing aggregation on the Adapter output but rather aggregates the Adapter at the parameter level. We adopt the same setting as the MoE methods, and the results could be calculated in the following way:
(19) $\Theta_{SMEAR}=\sum_{i=i}^{k}R(x)_{i}\Theta_{i},$
where $\Theta_{i}$ denote the parameter of the LoRA-$i$.
AdapterSoup AdapterSoup  (Chronopoulou et al., 2023), for new downstream tasks, retrieves the parameters that need to be involved in aggregation through sentence bert and performs weight-space averaging on these parameters to adapt to the new domain. We have uniformly retrieved 3 LoRAs for mixed-task to test their capabilities under mixed-task conditions.
LoRAHub LoRAHub  (Huang et al., 2023) also aggregates 20 LoRAs randomly for new downstream tasks. In order to learn the weight of LoRA, a black-box optimization method is employed to learn the weight of each LoRA without calculating the gradients of the large model. It performs weighted averaging at the parameter level. Similar to the training process of MoE, we randomly selected 20 samples for each task to form a unified training dataset for black-box optimization.
5.0.6. Implementation of RAMoLE.
To train RAMoLE, we divide our training process into two stages: first, training the LoraRetriever, and second, training the RouterLoRA. To train the LoraRetriever, we continue to perform instruction fine-tuning based on Instructor-xl  (Su et al., 2022). The training data consisted of only 40% of the tasks used to train task-specific LoRAs, with each task represented by 20 samples randomly selected from its respective LoRA training set. In this process, we categorized samples from the same LoRA as positive examples, while those from different LoRAs were considered negative examples.
5.0.7. Metrics.
Following Wei et al. (2021), we assess the performance on the ”Struct to Text” task using Rouge-{1, 2, L} and on the ”Translation” tasks using BLEU. Additionally, for the NLU tasks, we evaluate the exact match accuracy of each method.
Figure 4. DAMoLE Routing Distribution for NQ (Closed-book QA) and WMT-16 De/En (Translation). Red square brackets denote LoRAs from the same task cluster and yellow square brackets highlight the ideal LoRA for the respective tasks.
Figure 5. LoRA embedding similarity heatmap. Tasks from the same domain are grouped in square brackets.
5.1. Main Results
The main results of the mixed-task evaluation are shown in Tab.1. We present the mean performance across each task cluster. From the results, we have the following observations:
The full results are shown in Tab.2&3.
Figure 6. Comparison of Different LoRA Composition Strategies. The overall performance averages the IID and OOD results to demonstrate the comprehensive effectiveness of various LoRA composition strategies.
5.2. Analysis
Performance of Retriever.
We compare LoraRetriever with some popular off-the-shelf sentence embedding models in Huggingface and adopt the model nomenclature following Wolf et al. (2020). To analyze the effect of the percentage of tasks for training LoraRetriever, we trained three variants of LoraRetriever with different percentages. Tab.4 shows the performance of different retrieval models for retrieving relevant LoRAs. It is shown that guiding sentence embedding models with specific prompts leads to a performance improvement in retrieval compared to common retrieval models. After instruction fine-tuning, the retriever significantly enhanced the ability to retrieve the corresponding LoRA based on the input. Conducting instruction fine-tuning on 40% of the tasks resulted in a 2.36% increase in top-1 accuracy and a 9.80% increase in top-3 accuracy. Training across all tasks achieved the largest improvement. To demonstrate the generalizability of the proposed framework when dealing with unseen LoRAs, we used a retriever trained on 40% of the tasks in the main experiment to simulate the scenario of dynamic updates to the LoRA pool that might occur while providing services with LoraRetriever.
Performance of the On-the-fly MoLE Mechanism
Fig. 4 illustrates the routing distribution performance of the proposed mechanism across two tasks: Closed Book QA and Translation. The figure uses red square brackets to denote LoRAs from the same task cluster and yellow square brackets to highlight the ideal LoRA for the respective tasks. The results demonstrate that the proposed MoLE mechanism effectively assigns a higher weight to the ideal LoRA for each input at the module level. Additionally, the importance of higher-level layers in the routing process is evident, corroborating the findings of Gao et al. (2024).
Figure 7. Ablation of LoRA Dropout Strategy during RouterLoRA training.
Impact of the number of Retrieved LoRA
Effectiveness of Batch Inference Strategy.
To evaluate the efficiency of our proposed batch inference strategy, we compared the throughput of different batch sizes. Throughput is defined as the number of both input and output tokens per second across all requests in the mixed-task benchmark. We specifically compared the computational efficiency with that of a single LoRA. Our evaluation encompassed the entire evaluation dataset, and we limited the generation to the first produced token to mitigate discrepancies caused by varying generation lengths across different methods. These experiments were carried out on an NVIDIA A100 GPU (80GB) utilizing bfloat16 precision. As illustrated in Fig.8 (b), our batch inference strategy markedly improves the throughput of the framework, with a slight throughput reduction compared to a single LoRA. Notably, the Fusion strategy outperforms the mixture strategy in throughput efficiency, attributed to its parameter averaging approach, which circumvents the need for parallel computation across multiple LoRAs. The RAMoLE outperforms other composition methods by first gathering the corresponding LoRA for each input, allowing for a more effective inference process. As the batch size increases, the throughput of RAMoLE gradually reaches parity with that of a single LoRA.
Figure 8. The left figure shows the performance of LoraRetriever varying the number of LoRAs. The right figure shows the performance of Throughput varying the batch size.
Showcases.
Figure 9. Showcasing How the RAMoLE Framework Employs Multiple LoRAs for Cooperative Problem Solving.
6. Limitations and Potential Future Directions
7. Conclusion
Ethics Statement
In this work, we take the open-sourced Llama-2 (Touvron et al., 2023) model and Flan-v2 (Wei et al., 2021) dataset in the experiment. One potential ethical concern may arise when applying the proposed framework to real-world scenarios, as the data uploaded by LoRA contributors may contain user privacy. However, we believe this can be addressed by proper anonymization before uploading. Therefore, we believe our work will not pose a severe ethical concern.

### v2

@@ V-1 Training And Evaluation Datasets
V-3 Evaluation Dataset.
@@
For constructing the evaluation dataset, we randomly chose 50 samples from the test set for each task used in training 48 LoRAs, subsequently mixing and shuffling these samples to form a unified dataset with 2400 data entries.
V-4 Baseline Methods.
We compared our method with the following baselines: (1) Mixture of Experts  [33, 22, 11, 21, 24]. Many works have considered coordinating different adapters through MoE, and here we explored three distinct variants: one employing a soft mixture of experts and the other utilizing discrete routing (top1 and top3).It is worth noting that in this paper, our primary focus is on post-hoc MoE training. This implies that the LoRA experts have been previously trained, and only the parameters of the router are trainable during the training of the MoE methods. (2) SMEAR  [23] introduces the concept of adaptive routing by performing a weighted average of different adapters’ parameters to utilize various experts effectively. For the MoE and SMEAR baselines, challenges arise in scaling due to training confined to a limited set of LoRAs. Consequently, we strategically selected a dedicated LoRA expert for each domain to specialize in router training. (3) AdapterSoup  [26] uniformly selects the corresponding LoRAs for the entire downstream task, which lacks the ability to provide personalized service for diverse requests. (4) LoRAHub  [25] enables black-box optimization to learn the weights of various LoRA parameters, thereby facilitating weighted parameter averaging for specific downstream tasks. In our implementation, we conformed to the default setting, which entails randomly selecting 20 LoRAs from the available LoRA pool and performing weighted parameter averaging. For the MoE, SMEAR, and LoRAHub approaches, we selected 20 data samples from the training datasets of all tasks to serve as their training data.
TABLE II: Comparison of Sentence Embedding Techniques in LoRA Retrieval: The notation LoraRetrieverksignifies that the model underwent supplementary training on $k$ percent of the tasks. The performance of the selected retriever model in the evaluation phase is highlighted in gray.
Method Top 1 Top 3 Top 5 Top 8
all-mpnet-base-v2 58.40 78.26 84.77 90.24
all-MiniLM-L6-v2 51.73 73.11 80.54 87.18
msmarco-distilbert-cos-v5 45.84 66.01 75.14 82.67
gtr-t5-xl 53.19 69.72 77.41 83.59
LoraRetriever 0% 60.80 79.29 85.57 91.58
LoraRetriever 40% 63.16 89.09 95.45 98.97
LoraRetriever 100% 74.08 97.37 99.15 99.82
V-5 Implementation of Baseline Methods.
$y=\sum_{i=i}^{k}R(x)_{i}E_{i}(x).$ (16)
We implied two variants of the MoE routing mechanism. (1) Dense Gating. Following  [22], the router network consists of a dense layer with trainable parameter $W_{g}$, and the gating score could be obtained through a softmax function by:
$s_{i}=R(x)_{i}=softmax(W_{g}^{T}x),$ (17)
(2) Sparse Gate. To maintain the sparsity while training, we leverage the Gumbel softmax trick as  [23, 53], where the router can be written as:
$\hat{R}(x)_{i}=\frac{(log(R(x)_{i})+g_{i})/\tau}{\sum_{i=1}^{k}exp((log(R(x)_{i})+g_{i})/\tau)}$ (18)
where $g_{i}\sim\text{Gumbel}(0,1)$ and $\tau$ is the temperature.
SMEAR SMEAR  [23] does not perform routing aggregation on the Adapter output but rather aggregates the Adapter at the parameter level. We adopt the same setting as the MoE methods, and the results could be calculated in the following way:
$\Theta_{SMEAR}=\sum_{i=i}^{k}R(x)_{i}\Theta_{i},$ (19)
where $\Theta_{i}$ denote the parameter of the LoRA-$i$.
AdapterSoup AdapterSoup  [26], for new downstream tasks, retrieves the parameters that need to be involved in aggregation through sentence bert and performs weight-space averaging on these parameters to adapt to the new domain. We have uniformly retrieved 3 LoRAs for mixed-task to test their capabilities under mixed-task conditions.
LoRAHub LoRAHub  [25] also aggregates 20 LoRAs randomly for new downstream tasks. In order to learn the weight of LoRA, a black-box optimization method is employed to learn the weight of each LoRA without calculating the gradients of the large model. It performs weighted averaging at the parameter level. Similar to the training process of MoE, we randomly selected 20 samples for each task to form a unified training dataset for black-box optimization.
V-6 Implementation of RAMoLE.
To train RAMoLE, we divide our training process into two stages: first, training the LoraRetriever, and second, training the RouterLoRA. To train the LoraRetriever, we continue to perform instruction fine-tuning based on Instructor-xl  [27]. The training data consisted of only 40% of the tasks used to train task-specific LoRAs, with each task represented by 20 samples randomly selected from its respective LoRA training set. In this process, we categorized samples from the same LoRA as positive examples, while those from different LoRAs were considered negative examples.
V-7 Metrics.
Following [5], we assess the performance on the ”Struct to Text” task using Rouge-{1, 2, L} and on the ”Translation” tasks using BLEU. Additionally, for the NLU tasks, we evaluate the exact match accuracy of each method.
Figure 3: DAMoLE Routing Distribution for NQ (Closed-book QA) and WMT-16 De/En (Translation). Red square brackets denote LoRAs from the same task cluster and yellow square brackets highlight the ideal LoRA for the respective tasks.
Figure 4: LoRA embedding similarity heatmap. Tasks from the same domain are grouped in square brackets.
V-A Main Results
The main results of the mixed-task evaluation are shown in Tab.I. We present the mean performance across each task cluster. From the results, we have the following observations:
V-B Analysis
Performance of Retriever.
We compare LoraRetriever with some popular off-the-shelf sentence embedding models in Huggingface and adopt the model nomenclature following [54]. To analyze the effect of the percentage of tasks for training LoraRetriever, we trained three variants of LoraRetriever with different percentages. Tab.II shows the performance of different retrieval models for retrieving relevant LoRAs. It is shown that guiding sentence embedding models with specific prompts leads to a performance improvement in retrieval compared to common retrieval models. After instruction fine-tuning, the retriever significantly enhanced the ability to retrieve the corresponding LoRA based on the input. Conducting instruction fine-tuning on 40% of the tasks resulted in a 2.36% increase in top-1 accuracy and a 9.80% increase in top-3 accuracy. Training across all tasks achieved the largest improvement. To demonstrate the generalizability of the proposed framework when dealing with unseen LoRAs, we used a retriever trained on 40% of the tasks in the main experiment to simulate the scenario of dynamic updates to the LoRA pool that might occur while providing services with LoraRetriever.
Figure 5: Comparison of Different LoRA Composition Strategies. The overall performance averages the IID and OOD results to demonstrate the comprehensive effectiveness of various LoRA composition strategies.
Performance of the On-the-fly MoLE Mechanism
Fig. 3 illustrates the routing distribution performance of the proposed mechanism across two tasks: Closed Book QA and Translation. The figure uses red square brackets to denote LoRAs from the same task cluster and yellow square brackets to highlight the ideal LoRA for the respective tasks. The results demonstrate that the proposed MoLE mechanism effectively assigns a higher weight to the ideal LoRA for each input at the module level. Furthermore, the importance of higher-level layers in the routing process is evident, which corroborates the findings of [55].
Figure 6: Ablation of LoRA Dropout Strategy during RouterLoRA training.
Impact of the number of Retrieved LoRA
Effectiveness of Batch Inference Strategy.
To evaluate the efficiency of our proposed batch inference strategy, we compared the throughput of different batch sizes. Throughput is defined as the number of both input and output tokens per second across all requests in the mixed-task benchmark. We specifically compared the computational efficiency with that of a single LoRA. Our evaluation encompassed the entire evaluation dataset, and we limited the generation to the first produced token to mitigate discrepancies caused by varying generation lengths across different methods. These experiments were carried out on an NVIDIA A100 GPU (80GB) utilizing bfloat16 precision. As illustrated in Fig.7 (b), our batch inference strategy markedly improves the throughput of the framework, with a slight throughput reduction compared to a single LoRA. Notably, the Fusion strategy outperforms the mixture strategy in throughput efficiency, attributed to its parameter averaging approach, which circumvents the need for parallel computation across multiple LoRAs. RAMoLE outperforms other composition methods by first gathering the corresponding LoRA for each input, allowing for a more effective inference process. As the batch size increases, the throughput of RAMoLE gradually reaches parity with that of a single LoRA.
Figure 7: The left figure shows the performance of LoraRetriever varying the number of LoRAs. The right figure shows the performance of Throughput varying the batch size.
Showcases.
Figure 8: Showcasing How the RAMoLE Framework Employs Multiple LoRAs for Cooperative Problem Solving.
VI Limitations and Potential Future Directions
VII Conclusion
Appendix A Full Results
The full results on the evaluation dataset are shown in Tab.III & Tab.IV.
TABLE III: Mixed Tasks evaluation on both NLU & NLG tasks. “OOD” indicates that during retrieval, we masked the corresponding task’s LoRA for testing generalization when facing unknown tasks.
@@ Appendix A Full Results
TABLE IV: Mixed Tasks evaluation on both NLU & NLG tasks. “OOD” indicates that during retrieval, we masked the corresponding task’s LoRA for testing generalization when facing unknown tasks.
@@
Figure 9: Datasets and task clusters used to train LoRAs and generate mixed-task evaluation set in this paper (NLU tasks in blue; NLG tasks in green).
Appendix B Details of Training and Evaluation datasets
We leverage a subset of flan-v2 datasets  [5] as shown in Fig.9 for LoRA expert training and mixed-task dataset generation. We summarize the details of the used datasets as follows: